# Preconditioning without a preconditioner using randomized block Krylov subspace methods

### v2

@@ Abstract
preconditioning, conjugate gradient, block Krylov subspace methods
{MSCcodes}
@@ 1 Introduction
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\mu\geq 0$ is a critical task across the computational sciences. Systems of the form Eq. 1 arise in a variety of settings, including the following:
Positive definite linear systems. When $\mu=0$, Eq. 1 is simply the task of solving a positive definite linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$, one of the most common problems in numerical linear algebra.
Ridge regression. Given a data matrix $\mathbf{Z}\in\mathbb{R}^{n\times d}$, a source term $\mathbf{f}\in\mathbb{R}^{n}$, and a regularization parameter $\mu\geq 0$, the ridge regression problem is
(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\|\mathbf{Z}\mathbf{x}-\mathbf{f}\|^{2}+\mu\|\mathbf{x}\|^{2}.$
A direct computation shows that the solution to Eq. 2 is also the solution to
@@
i.e. a system of the form Eq. 1 with $\mathbf{A}=\mathbf{Z}^{\mathsf{T}}\mathbf{Z}$ and $\mathbf{b}=\mathbf{Z}^{\mathsf{T}}\mathbf{f}$. In a number of applications we are interested in the whole regularization path; i.e. $\mathbf{x}=\mathbf{x}_{\mu}$ for all values $\mu\geq 0$; e.g. in order to perform cross-validation to select a value of $\mu$ to use for future predictions.
@@
The term $(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}$ in the integrand of Eq. 4 is of the form Eq. 1 with $\mu=z^{2}$. Often we are interested in sampling several Gaussian vectors with the same covariance matrix.
Figure 1: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ in terms of matrix-loads (left) and wall-clock time (right) for our proposed randomized variant of the block conjugate gradient method (Refer to caption), standard conjugate gradient (Refer to caption), Nyström preconditioned conjugate gradient from [23] (Refer to caption), and generalizations Nyström preconditioned conjugate gradient using higher-depth Nyström approximations (Refer to caption). Our method outperforms all these methods without the need for selecting hyperparameters (see Theorem 3.4), which may be difficult to do effectively in practice. In this example, the runtime is dominated by the cost of loading chunks of the matrix into memory, so the wall-clock-time is nearly proportional to matrix-loads. Full experiment description in Section 7.1.
@@
Central to this paper are a class of algorithms for solving systems of the form Eq. 1 called Krylov Subspace Methods (KSMs). KSMs make use of the so-called Krylov subspace
(5) $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{t-1}\mathbf{b}\},$
which can be efficiently computed using matrix-vector products with $\mathbf{A}$. Since it is often of interest to solve Eq. 1 for multiple values of $\mu$, we repeatedly make use of the fact that the Krylov subspace is shift-invariant; i.e. that for any scalar $\mu$, $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})=\mathcal{K}_{t}(\mathbf{A},\mathbf{b})$.
@@
That is, $\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ is the space consisting of all linear combinations of vectors in the Krylov subspaces $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)}),\ldots,\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)})$.
As we discuss in Section 1.2, compared to their single-vector counterparts, block-KSMs have a number of computational benefits on modern computational architectures. As such, block-KSMs are widely used for tasks in numerical linear algebra such as low-rank approximation and eigenvalue approximation, where they are known to satisfy strong theoretical guarantees [27, 55, 34, 45, 38, 40, 61].
Interestingly, the use of block-KSMs for solving problems like Eq. 1 is more limited (see Section 5 for a discussion), and single-vector methods such as the conjugate gradient algorithm are more popular. A high-level aim of this paper is to encourage further study on the use of block-KSMs for tasks such as Eq. 1. In particular, we believe that techniques from randomized numerical linear algebra have the potential to provide new insights and theoretical guarantees for block-KSMs used to solve Eq. 1.
@@ 1.2 Computational assumptions
Throughout this paper, we assume $\mathbf{A}$ is accessed only through matrix-vector products $\mathbf{x}\mapsto\mathbf{A}\mathbf{x}$ or block matrix-vector products $\mathbf{x}_{1}~\cdots~\mathbf{x}_{m}]\mapsto[\mathbf{A}\mathbf{x}_{1}~\cdots~\mathbf{A}\mathbf{x}_{m}]$. While one can simulate a block matrix-vector product using $m$ matrix-vector products, in many settings (see e.g. Fig. 1) the cost of block products is often nearly independent of the block size $m$ (so long as $m$ is not too large). Allowing for more efficient data-access is widely understood as one of the major benefits of many randomized linear algebra algorithms [40, 61].
The algorithms we propose in this paper are designed to economize the number of times $\mathbf{A}$ is loaded into memory (henceforth referred to as matrix-loads), and not other costs such as matrix-vector products, floating point operations, or storage. In settings where matrix-loads are not the dominant cost, the proposed algorithms may not yield significant advantages. The costs of the algorithm from this paper are discussed more in Section 3.2, and we explore them in the numerical experiments in Section 7. A detailed understanding of tradeoffs in various costs is somewhat beyond the scope of this paper, but is of practical importance.
@@ 1.4 Organization
In Section 2 we describe key background on Krylov subspace methods and preconditioning. Our algorithm and main conceptual results are described in Section 3. Explicit and simple to use probabilistic results are proved in Section 4. We discuss our bounds in the context of past work in Section 5. Numerical experiments are provided in Section 7. Finally, we show how the theory we develop can be used to sample Gaussian vectors in Section 6.
@@ 2 Background
In this section we provide an overview of a number of relevant algorithms. The methods we discuss in this section are standard, and for mathematical simplicity, we define them by their optimality conditions rather than algorithmically (all of our bounds assume exact arithmetic); eee [29, 56, 42, etc.] for comprehensive treatments of these methods.
@@ 2.1 Preconditioned Conjugate Gradient
The preconditioned conjugate gradient algorithm (PCG) is a powerful KSM for solving positive definite linear systems. PCG makes use of a positive definite preconditioner $\mathbf{P}_{\mu}$ that transforms the system $\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b}$ into the system
(7) $(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\mathbf{y}=\mathbf{P}_{\mu}^{-1/2}\mathbf{b},\quad\mathbf{P}_{\mu}^{1/2}\mathbf{x}=\mathbf{y},$
@@
(8) $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu}):=\mathcal{K}_{t}(\mathbf{P}_{\mu}^{-1}\mathbf{A},\mathbf{P}_{\mu}^{-1}\mathbf{b}).$
@@
$\mathsf{cg}_{t}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A},\mathbf{b})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$
@@ 2.1.1 Convergence
PCG (and hence CG) satisfy a well-known convergence guarantee in terms of the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$; see e.g. [29]. [^1] Here we use that $(({\sqrt{\kappa}-1})/({\sqrt{\kappa}+1}))^{t}\leq\exp(-2t/\sqrt{\kappa})$ for all $\kappa\geq 1$ in order to obtain a simpler expression than the typical bound. This approximation is pessimistic when $\kappa\approx 1$, but it’s clear from our analysis that other bounds for CG can be used instead.
@@
Corollary 2.3 implies that if we can find $\mathbf{P}_{\mu}$ so that $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is small, then PCG converges rapidly. The choice of $\mathbf{P}_{\mu}$ minimizing this condition number is of course $\mathbf{P}_{\mu}=\mathbf{A}_{\mu}$, but this is not a practical preconditioner; if we knew $\mathbf{A}_{\mu}^{-1}$ we could easily compute the solution $\mathbf{A}_{\mu}^{-1}\mathbf{b}$. Thus, finding a suitable choice of $\mathbf{P}_{\mu}$, which balances improvements to the convergence of PCG with the cost of building/applying $\mathbf{P}_{\mu}$ is critical.
@@ 2.2 Deflation preconditioner
(10) $\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
This means applying $\mathbf{P}_{\mu}^{-1}$ to a vector can be done in $O(dr)$ arithmetic operations, which is relatively cheap if $r\ll d$. Hence, so long as a reasonable factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ can be obtained, Eq. 9 can be used as a preconditioner; see [24, 33] and the references within.
@@ 2.2.1 Exact deflation
Traditionally, it has been suggested to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as $\mathbf{A}]\!]_{r}$, the best rank-$r$ approximation to $\mathbf{A}$; i.e. to set $\mathbf{D}$ as the top $r$ eigenvalues of $\mathbf{A}$ and $\mathbf{U}$ the corresponding eigenvectors. In this case Eq. 9 is sometimes called a spectral deflation preconditioner. The name arises because the eigenvalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ are
(11) $\underbrace{\theta+\mu,\ldots,\theta+\mu}_{\text{$r$-copies}},\lambda_{r+1}+\mu,\ldots,\lambda_{d}+\mu,$
and hence, if $\theta\in[\lambda_{d},\lambda_{r+1}]$, then the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ is $\kappa_{r+1}(\mu)=(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)$. In other words, the top $r$ eigenvalues of $\mathbf{A}_{\mu}$ are deflated. In combination with Corollary 2.3, this yields the following convergence guarantee.
@@
Let $\mathbf{P}_{\mu}$ be the preconditioner Eq. 9 corresponding to $\mathbf{A}]\!]_{r}$, the rank-$r$ truncated SVD of $\mathbf{A}$, and $\theta\in[\lambda_{d},\lambda_{r+1}]$. Then the $t$-th PCG iterate (Definition 2.1) corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
@@
When $\lambda_{r+1}\ll\lambda_{1}$ and $\mu$ is not too large, then the rate of convergence guaranteed by the bound can be much faster than without preconditioning (r=0r=0). [^2] Standard CG also satisfies bounds in terms of $\lambda_{r+1}/\lambda_{d}$, at least in exact arithmetic; see Theorem B.1 and the discussion in Appendix B. Of course, any potential benefits to convergence must be weighted against the cost of constructing the spectral deflation preconditioner, and exact deflation, which requires computing the top eigenvectors of $\mathbf{A}$, can be costly.

### v3

@@ Abstract
preconditioning, randomized, conjugate gradient, block Krylov subspace methods
{AMS}
@@ 1 Introduction
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\mu\geq 0$ is a critical task across the computational sciences. Systems of the form Equation˜1 arise in a variety of settings, including the following:
Positive definite linear systems. When $\mu=0$, Equation˜1 is simply the task of solving a positive definite linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$, one of the most common problems in numerical linear algebra.
Ridge regression. Given a data matrix $\mathbf{Z}\in\mathbb{R}^{n\times d}$, a source term $\mathbf{f}\in\mathbb{R}^{n}$, and a regularization parameter $\mu\geq 0$, the ridge regression problem***Ridge regression is a special case of Tikhonov regularization, where $\mu\|\mathbf{x}\|^{2}$ is replaced by a somewhat more general regularization term $\|\mathbf{\Gamma}\mathbf{x}\|^{2}$ [22]. is
(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\left(\|\mathbf{Z}\mathbf{x}-\mathbf{f}\|^{2}+\mu\|\mathbf{x}\|^{2}\right).$
A direct computation shows that the solution to Equation˜2 is also the solution to
@@
i.e. a system of the form Equation˜1 with $\mathbf{A}=\mathbf{Z}^{\mathsf{T}}\mathbf{Z}$ and $\mathbf{b}=\mathbf{Z}^{\mathsf{T}}\mathbf{f}$. In a number of applications we are interested in the whole regularization path; i.e. $\mathbf{x}=\mathbf{x}_{\mu}$ for all values $\mu\geq 0$; e.g. in order to perform cross-validation to select a value of $\mu$ to use for future predictions.
@@
The term $(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}$ in the integrand of Equation˜4 is of the form Equation˜1 with $\mu=z^{2}$. Often we are interested in sampling several Gaussian vectors with the same covariance matrix.
Figure 1: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ in terms of matrix-loads (left) and wall-clock time (right) for our proposed randomized variant of the block conjugate gradient method (Refer to caption), standard conjugate gradient (Refer to caption), Nyström preconditioned conjugate gradient from [25] (Refer to caption), and generalizations Nyström preconditioned conjugate gradient using higher-depth Nyström approximations (Refer to caption). Our method outperforms all these methods without the need for selecting hyperparameters (see Theorem˜3.4), which may be difficult to do effectively in practice. In particular, we store $\mathbf{A}$ in 8 separate $1000\times 8000$ chunks and perform (block) matrix-vector products with $\mathbf{A}$ by sequentially loading a single chunk from the disk into random access memory and performing the appropriate part of the products. The runtime is dominated by the cost of loading chunks of the matrix into memory, so the wall-clock-time is nearly proportional to matrix-loads. Full experiment description in Subsection˜7.1.
@@
All three problems described above are commonly solved using a class of algorithms called Krylov Subspace Methods (KSMs); see e.g., [63, 39, 8]. KSMs make use of the so-called Krylov subspace
$\mathcal{K}_{t}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{t-1}\mathbf{b}\},$
which can be efficiently computed using matrix-vector products with $\mathbf{A}$. Since it is often of interest to solve Equation˜1 for multiple values of $\mu$, we repeatedly make use of the fact that the Krylov subspace is shift-invariant; i.e. that
(5) $\forall\mu\in\mathbb{C}:\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})=\mathcal{K}_{t}(\mathbf{A},\mathbf{b}).$
@@
That is, $\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ is the space consisting of all linear combinations of vectors in the Krylov subspaces $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)}),\ldots,\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)})$. Analogous to Equation˜5, the block Krylov subspace is shift invariant; i.e. $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ for any shift $\mu$.
As we discuss in Subsection˜1.2, compared to their single-vector counterparts, block-KSMs have a number of computational benefits on modern computational architectures. As such, block-KSMs are widely used for tasks in numerical linear algebra such as low-rank approximation and eigenvalue approximation, where they are known to satisfy strong theoretical guarantees [31, 62, 38, 51, 43, 46, 68].
Interestingly, the use of block-KSMs for solving problems like Equation˜1 is more limited (see Section˜6 for a discussion), and single-vector methods such as the conjugate gradient algorithm are more popular. A high-level aim of this paper is to encourage further study on the use of block-KSMs for tasks such as Equation˜1. In particular, we believe that techniques from randomized numerical linear algebra have the potential to provide new insights and theoretical guarantees for block-KSMs used to solve Equation˜1.
@@ 1.2 Computational assumptions
Throughout this paper, we assume $\mathbf{A}$ is accessed only through matrix-vector products $\mathbf{x}\mapsto\mathbf{A}\mathbf{x}$ or block matrix-vector products $\mathbf{x}_{1}~\cdots~\mathbf{x}_{m}]\mapsto[\mathbf{A}\mathbf{x}_{1}~\cdots~\mathbf{A}\mathbf{x}_{m}]$. While one can simulate a block matrix-vector product using $m$ matrix-vector products, in many settings the cost of block products is often nearly independent of the block size $m$ (so long as $m$ is not too large). For example, this can be the case when $\mathbf{A}$ is so large that it must be loaded in chunks from slow memory (as in Figure˜1), where $\mathbf{A}$ is the Hessian of a large neural-network and matrix-vector products are performed using “Pearlmutter’s trick” [58], or where $\mathbf{A}$ corresponds to the solution operator of an integral equation and is applied via a solver fast direct solver [45]. Allowing for more efficient data-access is widely understood as one of the major benefits of many randomized linear algebra algorithms [46, 68].
The algorithms we propose in this paper are designed to economize the number of times $\mathbf{A}$ is loaded into memory (henceforth referred to as matrix-loads), and not other costs such as matrix-vector products, floating point operations, or storage. In settings where matrix-loads are not the dominant cost, the proposed algorithms may not yield significant advantages. The costs of the algorithm from this paper are discussed more in Subsection˜3.2, and we explore them in the numerical experiments in Section˜7. A detailed understanding of tradeoffs in various costs is somewhat beyond the scope of this paper, but is of practical importance.
@@ 1.4 Organization
In Section˜2 we describe key background on Krylov subspace methods and preconditioning. Our algorithm and main conceptual results are described in Section˜3. Explicit and simple to use probabilistic results are proved in Section˜4. We discuss our bounds in the context of past work in Section˜6 and show how the theory we develop can be used to sample Gaussian vectors in Section˜5. Finally, numerical experiments are provided in Section˜7.
@@ 2 Background
In this section we provide an overview of a number of relevant algorithms. The methods we discuss in this section are standard, and for mathematical simplicity, we define them by their optimality conditions rather than algorithmically (all of our bounds assume exact arithmetic); see e.g., [33, 63, 48] for comprehensive treatments of these methods.
@@ 2.1 Preconditioned Conjugate Gradient
The preconditioned conjugate gradient algorithm (PCG) is a powerful KSM for solving positive definite linear systems. PCG makes use of a symmetric positive definite preconditioner $\mathbf{P}_{\mu}$ that transforms the system $\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b}$ into the system
$(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\mathbf{y}=\mathbf{P}_{\mu}^{-1/2}\mathbf{b},\quad\mathbf{P}_{\mu}^{1/2}\mathbf{x}=\mathbf{y},$
@@
(7) $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu}):=\mathcal{K}_{t}(\mathbf{P}_{\mu}^{-1}\mathbf{A}_{\mu},\mathbf{P}_{\mu}^{-1}\mathbf{b}).$
@@
$\mathsf{cg}_{t}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$
@@ 2.1.1 Convergence
PCG (and hence CG) satisfy a well-known convergence guarantee in terms of the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$; see e.g. [33].Here we use that $(({\sqrt{\kappa}-1})/({\sqrt{\kappa}+1}))^{t}\leq\exp(-2t/\sqrt{\kappa})$ for all $\kappa\geq 1$ in order to obtain a simpler expression than the typical bound. This approximation is pessimistic when $\kappa\approx 1$, but it’s clear from our analysis that other bounds for CG can be used instead.
@@
Corollary˜2.3 implies that if we can find $\mathbf{P}_{\mu}$ so that $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is small, then PCG converges rapidly. The choice of $\mathbf{P}_{\mu}$ minimizing this condition number is of course $\mathbf{P}_{\mu}=\mathbf{A}_{\mu}$, but this is not a practical preconditioner; if we knew $\mathbf{A}_{\mu}^{-1}$ we could easily compute the solution $\mathbf{A}_{\mu}^{-1}\mathbf{b}$. Thus, finding a suitable choice of $\mathbf{P}_{\mu}$, which balances improvements to the convergence of PCG with the cost of building/applying $\mathbf{P}_{\mu}$ is critical.
@@ 2.2 Deflation preconditioner
$\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
This means applying $\mathbf{P}_{\mu}^{-1}$ to a vector can be done in $O(dr)$ arithmetic operations, which is relatively cheap if $r\ll d$. Hence, so long as a reasonable factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ can be obtained, Equation˜8 can be used as a preconditioner; see [26, 37] and the references within.
@@ 2.2.1 Exact deflation
Traditionally, it has been suggested to take $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ as $\mathbf{A}]\!]_{r}$, the best rank-$r$ approximation to $\mathbf{A}$; i.e. to set $\mathbf{D}$ as a diagonal matrix containing the top $r$ eigenvalues of $\mathbf{A}$ and $\mathbf{U}$ the corresponding eigenvectors. In this case Equation˜8 is sometimes called a spectral deflation preconditioner. The name arises because the eigenvalues of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ are
$\underbrace{\theta+\mu,\ldots,\theta+\mu}_{\text{$r$-copies}},\lambda_{r+1}+\mu,\ldots,\lambda_{d}+\mu,$
and hence, if $\theta\in[\lambda_{d},\lambda_{r+1}]$, then the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ is $\kappa_{r+1}(\mu)=(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)$. In other words, the top $r$ eigenvalues of $\mathbf{A}_{\mu}$ are deflated. In combination with Corollary˜2.3, this yields the following convergence guarantee.
@@
Let $\mathbf{P}_{\mu}$ be the preconditioner Equation˜8 corresponding to $\mathbf{A}]\!]_{r}$, the rank-$r$ truncated SVD of $\mathbf{A}$, and $\theta\in[\lambda_{d},\lambda_{r+1}]$. Then the $t$-th PCG iterate (Definition˜2.1) corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
@@
When $\lambda_{r+1}\ll\lambda_{1}$ and $\mu$ is not too large, then the rate of convergence guaranteed by the bound can be much faster than without preconditioning ($r=0$).Standard CG also satisfies bounds in terms of $\lambda_{r+1}/\lambda_{d}$, at least in exact arithmetic; see B.1 and the discussion in B. Of course, any potential benefits to convergence must be weighted against the cost of constructing the spectral deflation preconditioner, and exact deflation, which requires computing the top eigenvectors of $\mathbf{A}$, can be costly.